# Contextual Evaluation of Large Language Models for Classifying Tropical and Infectious Diseases

### v1

@@ 1 Introduction
Neglected tropical diseases, while extremely preventable and treatable, continue to be highly prevalent in the poorest regions of the world, affecting 1.7 billion people globally with disproportionate impacts on women and children [1]. Similarly, half of the world’s population are at risk from infectious diseases, which continue to lead in global mortality and morbidity, resulting in more than 52 million deaths globally, with 99% of diseases occurring in developing countries [2]. Challenges in preventing and treating these diseases include ongoing tracking, early detection, accurate initial diagnosis, management and vaccine limitations [1, 3]. During the pandemic various scalable measures were implemented to address these challenges specific to COVID-19 [4, 5, 6, 7, 8]. Other tropical and infectious diseases have had limited attention for tracking and surfacing accurate diagnosis. The use of large language models for health related question-answering has increased in recent years demonstrating applications for a variety of health use cases [9, 10, 11, 12]. However, there is limited work that has focused on tropical and infectious diseases. These are particularly of interest as they may present out-of-distribution cases, given that they mostly occur in the global south which may be underrepresented in training and evaluation datasets and may lead to potential biases [13, 14]. Additionally there is limited understanding of how different contextual factors such as demographics, prompt styles, and subsets of information (eg. symptoms only, versus symptoms+location) may influence model performance.
Few papers assess the use of machine learning and LLMs for tropical and infectious diseases. Shenoy et al. 2022 [15] found in a study of 40 tropical and infectious disease clinicians, 35 indicated the need for a decision making tool. They compare logistic regression and decision trees for binary classification and find an average prediction accuracy of 79-84%. Pfohl et al. 2024 [16] developed a health equity dataset which introduced LLM evaluation of a tropical and infectious diseases (TRINDs) dataset and had clinicians and health equity experts evaluate LLM responses to tropical and infectious disease related questions,finding evidence of biased responses. Mondal et al. 2024 [17] developed 50 simulated infectious disease cases with histories, lab reports, imaging findings and evaluated it on 4 different general LLMs and compared the quantity of differential diagnosis to that of medical experts. They found that LLMs generally had difficulties matching the experts’ differential diagnosis. Schwartz et al. 2024 [18] delineate failure points for LLMs for infectious diseases consultation to clinical workflow questions and find that LLMs provide dangerous hallucinations and harmful advice for disease management-type questions.
In this paper we build on the TRopical and INfectious Diseases (TRINDs) dataset [16]. We make the following primary contributions:
@@ 2 Methods
Original TRINDs dataset Generalist LLM vs specialist LLM accuracy
Comparison of LLM and human expert performance
@@
Multiple choice set Comparison of LLM and human expert performance
@@ 2.1 Dataset generation and expansion:
Original TRINDs dataset: We base this work on the original TRINDs dataset [16]. To summarize, the authors examined authoritative sources containing factual information about different diseases to create a dataset of 106 questions pertaining to tropical and infectious diseases across different regions around the world. We use a subset of 52 questions from the TRINDs data as a seed set. Each question in the seed set follows a templated structure where a patient persona is presented with general symptoms, direct attributes, specific symptoms, as well as context and lifestyle/risk factors. Each question is associated with a ground truth disease label, which were reviewed by clinicians for accuracy.
@@
Counterfactual sets: Location: We examined how location influences LLMs responses. The original dataset was created with locations where the disease had a known likelihood of occurring. For each of the original dataset and contextual subsets that included a location, we switched out the original location for a single counterfactual location where there was less probability of disease occurring. Here we used “San Francisco” as that location to generate 52 counterfactual location queries. Race: We created versions of the original dataset where we included a race input (Asian, Black and White) for each disease, yielding 159 additional queries. Gender: We created versions of the dataset that had only male, only female and non-binary demographics, yielding an additional 159 queries.
French language set: Given prevalence of tropical and infectious diseases in non-English speaking countries we sought to understand how language influenced the performance of the model. A researcher whose official language was French manually translated the original dataset of 52 queries and diagnosis into French and compared performance to the English dataset. We selected French as our primary interest lies in the African continent and official languages that are used are primarily in English and French. This led to 52 prompts.
Multiple choice set: To compare the LLMs to expert baselines, we generate 153 multiple choice questions from the original, expanded demographics and expanded semantics sets. We input the questions and ground truth diagnosis from the original dataset and prompted an LLM to provide 5 multiple choice options that included the ground truth and 4 other broad tropical and infectious disease options.
@@
LLM-Expanded semantic set: As prompts from users may present in a variety of ways, and as symptoms may vary as well, we synthetically expanded the dataset to 2651 additional queries. These were created by instructing the LLM to first generate a demographically altered version of the question, in the same way as described above for the expanded demographics set, and to then follow that up by another alteration, generating an alternative formulation of the question for the given patient and disease, but describing a different set of common symptoms related to the disease, and potentially a different personal patient background of relevance for disease risk or onset. LLM-based filtering was applied to these generations as well, where proposals would get discarded in case the generated symptoms did not match the disease, or the proposed demographic does not appear to be plausible.
@@
In total we generated a dataset of 11,719 queries from the original seed set of 52 queries across 50 tropical and infectious diseases. Refer to Appendix A.2 for examples.
@@ 2.2 Model evaluation
We use two baseline models- Gemini Ultra [19], a generalist large language model (standard hyperparameters: batchsize 16, temperature 0.7, top_k=32), and MedLM Medium [20], a LLM specialized for the health domain (standard hyperparameters: batchsize 32, temperature 0.2, top_k=None). As a baseline, we prompt-tuned both models providing instructions and 2-shot examples to guide the model’s output (Appendix A.3). The models were prompted to provide an output of the ground truth disease label. This was repeated 4 additional times to yield a total of 5 outcomes per experiment for statistical analysis.
@@ 2.3 Auto-rater LLM Evaluations
We developed an automated rater to score each query for accuracy out of 5 repeated runs. This was developed by prompting an LLM to determine whether the words were either structural or meaningfully similar to each other and score them as correct if they were. For instance if the ground truth was Taeniasis/cysticercosis and model output was Tapeworm it would be marked as correct, since these are meaningfully similar, even if they are structurally different.We performed a manual review of the automated rating on a subset of the data to optimize the rating process and ensure there were no errors.
@@ 2.4 Human Expert Baseline
Table 2: Expert panel demographics (n=7)
@@
We created a human expert baseline study to understand expert performance on the dataset. The study was deemed IRB exempt by an internal ethics review personnel. We recruited experts- public health researchers, and medical doctors- who had generalizable knowledge of TRINDs, using a snowballing approach via our networks.
@@ 2.6 Statistical analysis
We performed t-test statistical analysis to compare performance between medical specific and general purpose models for each experiment, original and counterfactual datasets, base and many-shot prompt tuned models, and LLM and human expert performance.
@@ 3 Results
Figure 1: Model performance on persona variations. a) Generalist (Gemini) and specialist (MedLM) model performance on clinical, consumer and French persona variations. b) Gemini model performance on counterfactual location inputs. c) Gemini performance for contextual combinations of attributes and factors and count. d) Gemini performance for race counterfactuals e) Gemini performance for gender counterfactuals. Error bars are 90% confidence interval. *=p<0.05
@@ 3.1.2 Assessing Gemini model performance on varied combinations of attributes and factors
We found that symptoms, location and risk factors enable the best model performance, followed by the full persona of symptoms, location, risk factors and personal attributes (Figure 1c). This implies that excluding personal attributes such as age and gender, may preserve privacy while still maintaining beneficence. The worst performing contextual combinations were symptoms alone (46.8%), symptoms and attribute (44.91%) and general symptoms, location, attributes and risk (44.5%), demonstrating the need for both specific and general symptoms, as well as other contextual factors such as location and risk factors to enhance model accuracy. We also show performance per disease and find that the model performs best on relatively widespread or distinctive diseases eg. Trachoma, HIV and Rabies, but performs worst on less common or diseases with common symptoms eg. tuberculosis (Figure 2a).
@@ 3.1.3 Assessing Gemini model performance on counterfactual inputs
Location: We found that generally, the counterfactual location decreased performance across all contextual combinations that included location, but performance was least affected in cases when full context persona, or combination of symptom, location and risk demonstrating the usefulness of including added contextual information (Figure 1b). The location counterfactual caused the model to perform the worst when only symptoms were provided or when only general symptoms, or specific symptoms were provided with other information, demonstrating the need to have both specific and general symptoms, plus other contextual information. We find similar trends as well across individual diseases (Figure 2b). Race: There were no statistically significant differences in performance across different racial counterfactual inputs, even though we note that performance for “White race” counterfactuals increased slightly (Figure 1d). Gender: There were no statistically significant differences in performance across different gender counterfactual inputs, even though we note that performance for each of the gender counterfactuals was slightly higher than the original persona (Figure 1e).
@@ 3.1.4 Assessing model performance on the demographic and semantic expansions
Figure 3: Model performance on expanded dataset. a) LLM performance on demographic clinical and consumer augmentations (2635 eacg) and b)LLM performance on semantic clinical and consumer augmentations (2651 each). We compared the base model, with the multi-shot tuned model. ****=p<0.00005
@@ 3.2.2 Expert data rating and qualitative feedback:
Experts rated level of helpfulness of each of the contextual information in providing a diagnosis. Symptoms were rated highest by most experts, followed by risk factor, location and then demographics (Figure 4 c). This is in line with LLM results demonstrating Symptoms, location and risk factors as being most important for classification. On the dataset quality, experts generally indicated high scores (4-5) for accuracy, completeness, range of tropical and infectious diseases covered, geographic and demographic diversity, and timelines of the dataset (Figure 4 b). Experts rated the level of difficulty of the questions as 3-4/5. The diversity in the style of questions- asking was rated low with several experts commenting on the repetitiveness of the questions. Experts recommended including the use of over-the-counter medications, less repetition in the question style, improving the specificity of some of the symptoms, and the need for differential diagnosis. There were also comments on improving the quality of the LLM-generated queries. One expert commented on the inclusion of smallpox which has been eradicated for 40 years, but indicated the potential tie to the ongoing Monkeypox (Mpox) epidemic. Experts also indicated the need for patient images to provide more informative responses, which we plan to leverage in a multimodal version of the dataset.
@@
4 Limitations and Future work
Limitations of this work include the focus on only disease classification, primary focus on English and primary focus on text-based queries. Future work could explore evaluating other tasks such as management steps and treatments, additional languages and multimodal datasets such as relevant disease-related images, or sounds from coughs and breathing.
@@ 5 Conclusion
Overall this work finds that while LLM performance on providing a diagnosis of tropical and infectious diseases is low, performance improves with simple in-context learning with our dataset. We find that larger generalist models outperform smaller specialist models. We underline the importance of context in performance, noting that providing symptoms, risk factors and location outperform the provision of symptoms alone to the LLM. This work provides a scalable methodology for evaluating LLMs for health in global settings for out-of-distribution cases.
6 Acknowledgements
We would like to acknowledge Dr. Lynda Osadebe for her expert insights on tropical and infectious diseases.

### v3

# neurips_2024.bib

@@ 1 Introduction
Neglected tropical diseases, while extremely preventable and treatable, continue to be highly prevalent in the poorest regions of the world, affecting 1.7 billion people globally with disproportionate impacts on women and children [hudu2024insight]. Similarly, half of the world’s population are at risk from infectious diseases, which continue to lead in global mortality and morbidity, resulting in more than 52 million deaths globally, with 99% of diseases occurring in developing countries [rana2023changes]. Challenges in preventing and treating these diseases include surveillance, early detection, accurate initial diagnosis, management and vaccine limitations [hudu2024insight, vuitika2022vaccines]. During the pandemic various scalable measures were implemented to address these challenges specific to COVID-19 [bragazzi2020big, gupta2020analysis, mbunge2020integrating, marbouh2020blockchain, lee2020global]. Other tropical and infectious diseases have had limited attention for surveillance and surfacing accurate diagnosis.
The use of large language models for health-related question-answering has increased in recent years demonstrating applications for a variety of health use cases [sallam2023chatgpt, tian2024opportunities, lu2024large, mannhardt2024impact]. However, there is limited work that has focused on tropical and infectious diseases. These are particularly of interest as they may present out-of-distribution cases, given that they mostly occur in the global south which may be underrepresented in training and evaluation datasets and may lead to potential biases [asiedu2023globalizing, asiedu2024case]. Additionally there is limited understanding of how different contextual factors such as demographics, prompt styles, and subsets of information (eg. symptoms only, versus symptoms+location) may influence model performance.
[…]
In this paper we build on the TRopical and INfectious Diseases (TRINDs) dataset [pfohl2024toolbox], making the following primary contributions:
@@ 2 Methods
Original TRINDs dataset Generalist LLM vs. specialist LLM accuracy
LLM vs. human expert performance
@@
Multiple choice set LLM vs. human expert performance
@@ 2.1 Dataset generation and expansion:
Original TRINDs dataset: We base this work on the original TRINDs dataset [pfohl2024toolbox]. To summarize, the authors examined authoritative sources containing factual information about different diseases to create a dataset of 106 questions pertaining to tropical and infectious diseases across different regions around the world. We use a subset of 52 questions from the TRINDs data as a seed set. Each question in the seed set follows a templated structure where a patient persona is presented with general symptoms, direct attributes, specific symptoms, as well as context and lifestyle/risk factors. Each question is associated with a ground truth disease label, which were reviewed by clinicians for accuracy.
@@
Counterfactual sets: Location: We examined how location influences LLM responses. The original dataset was created with locations where the disease had a known likelihood of occurring. For each of the original dataset and contextual subsets that included a location, we switched out the original location for a single counterfactual location where there was less probability of disease occurring. Here we used “San Francisco” as that location to generate 52 counterfactual location queries. Race: We created versions of the original dataset where we included a race input (Asian, Black and White) for each disease, yielding 159 additional queries. Gender: We created versions of the dataset that had only male, only female and non-binary demographics, yielding an additional 159 queries.
French language set: Given prevalence of tropical and infectious diseases in non-English speaking countries we sought to understand how language influenced the performance of the model. A researcher whose official language is French, manually translated the original dataset of 52 queries and diagnosis into French and compared performance to the English dataset. We selected French as our primary interest lies in the African continent and official languages that are used are primarily in English and French. This led to 52 prompts.
Multiple choice set: To compare the LLMs to expert baselines, we generate 153 multiple choice questions from the original, expanded demographics and expanded semantics sets. We input the questions and ground truth diagnosis from the original dataset and prompted an LLM to provide 5 multiple choice options that included the ground truth label, and 4 other broad tropical and infectious disease options.
@@
LLM-Expanded semantic set: As prompts from users may present in a variety of ways, and as symptoms may vary as well, we synthetically expanded the dataset, adding 2651 additional queries. These were created by instructing the LLM to first generate a demographically altered version of the question, in the same way as described above for the expanded demographics set. The LLM then follows that up by another alteration, generating an alternative formulation of the question for the given patient and disease, but describing a different set of common symptoms related to the disease, and potentially a different personal patient background of relevance for disease risk or onset. LLM-based filtering was applied to these generations as well, where proposals would get discarded in case the generated symptoms did not match the disease, or the proposed demographic does not appear to be plausible.
@@
In total we generated a dataset of 11,719 queries from the original seed set of 52 queries across 50 tropical and infectious diseases. We summarize the dataset and experiments run in Table 1. Refer to Appendix A.2 for examples.
@@ 2.2 Model evaluation
We use two baseline models- Gemini Ultra [gemini], a generalist large language model (standard hyperparameters: batchsize 16, temperature 0.7, top_k=32), and MedLM Medium [medlm], a LLM specialized for the health domain (standard hyperparameters: batchsize 32, temperature 0.2, top_k=None). As a baseline, we prompt-tuned both models, providing instructions and 2-shot examples to guide the model’s output (Appendix A.3). The models were prompted to provide an output of the ground truth disease label. This was repeated 4 additional times to yield a total of 5 outcomes per experiment for statistical analysis.
@@ 2.3 Auto-rater LLM Evaluations
We developed an automated rater to score each query for accuracy out of the 5 repeated runs. This was developed by prompting an LLM to determine whether the words were structurally and/or meaningfully similar to each other and to score them as correct if they were. For instance if the ground truth was Taeniasis/cysticercosis and model output was Tapeworm it would be marked as correct, since these are meaningfully similar, even if they are structurally different. We performed a manual review of the automated rating on a subset of the data to optimize the rating process and ensure there were no errors.
@@ 2.4 Human Expert Baseline
Table 2: Expert panel demographics (n=7).
@@
We created a human expert baseline study to understand expert performance on the dataset. The purpose of this was to determine how experts in tropical and infectious diseases performed on a representative sample of the data to enable human expert vs LLM comparison. The study was deemed IRB exempt by an internal ethics review personnel. We recruited experts- public health researchers, and medical doctors- who had generalizable knowledge of TRINDs, using a snowballing approach via our networks.
@@ 2.6 Statistical analysis
We performed two-tailed student’s t-test (alpha=0.025) for statistical analysis to compare performance between medical specific and general purpose models for each experiment, original and counterfactual datasets, base and many-shot prompt tuned models, and LLM and human expert performance.
@@ 3 Results
Figure 1: Model performance on persona variations. a) Generalist (Gemini) and specialist (MedLM) model performance on clinical, consumer and French persona variations. b) Gemini model performance on counterfactual location inputs. c) Gemini performance for contextual combinations of attributes and factors and count. d) Gemini performance for race counterfactuals e) Gemini performance for gender counterfactuals. Legend: S=symptoms(general and specific), gS=general symptoms, sS= specific symptoms, L=location, A=attribute (age and gender) Error bars are 90% confidence interval. *=p<0.025
@@ 3.1.2 Assessing Gemini model performance on varied combinations of attributes and factors
We found that symptoms, location and risk factors (SLR) enable the best model performance, followed by the full persona (FP) of symptoms, location, risk factors and personal attributes (Figure 1c). This implies that excluding personal attributes such as age and gender, may preserve privacy while still maintaining beneficence. The worst performing contextual combinations were general and specific symptoms (S) alone (46.8%), symptoms and attribute (SA) (44.91%) and general symptoms, location, attributes and risk factors (gSLAR) (44.5%), demonstrating the need for both specific and general symptoms, as well as other contextual factors such as location and risk factors to enhance model accuracy. We also show performance per disease and find that the model performs best on relatively widespread or distinctive diseases eg. Trachoma, HIV and Rabies, but performs worst on less common or diseases with common symptoms eg. tuberculosis (Figure 2a).
@@ 3.1.3 Assessing Gemini model performance on counterfactual inputs
Location: We found that generally, the counterfactual location decreased performance across all contextual combinations that included location (L), but performance was least affected in cases with full persona (FP), or with the combination of general and specific symptoms, location and risk (SLR) demonstrating the usefulness of including added contextual information (Figure 1b). Counterfactual location caused the model to perform the worst when only symptoms were provided without any additional context, or when either general symptoms or specific symptoms were provided even with other information. This demonstrates the need to have both specific and general symptoms, plus other contextual information for optimal performance. We find that this pattern is consistent across the individual diseases (Figure 2b). Race: There were no statistically significant differences in performance across different racial counterfactual inputs, even though we note that performance for “White race” counterfactuals increased slightly (Figure 1d). Gender: There were no statistically significant differences in performance across different gender counterfactual inputs, even though we note that performance for each of the gender counterfactuals was slightly higher than the original persona (Figure 1e).
@@ 3.1.4 Assessing model performance on the demographic and semantic expansions
Figure 3: Model performance on expanded dataset. a) LLM performance on demographic clinical and consumer augmentations (2635 each) and b)LLM performance on semantic clinical and consumer augmentations (2651 each). We compared the base model, with the multi-shot tuned model. ****=p<0.00005
@@ 3.2.2 Expert data rating and qualitative feedback:
Experts rated level of helpfulness of each of the contextual information in providing a diagnosis. Symptoms were rated highest by most experts, followed by risk factors, location and then demographics attributes (Figure 4 c). This is in line with LLM results that indicate that symptoms, location and risk factors are most important for LLM classification. On the dataset quality, experts generally indicated high scores (4-5) for accuracy, completeness, range of tropical and infectious diseases covered, geographic and demographic diversity, and timelines of the dataset (Figure 4 b). Experts rated the level of difficulty of the questions as 3-4/5. The diversity in the style of questions-asking was rated low with several experts commenting on the repetitiveness of the questions. Experts recommended including the use of over-the-counter medications, less repetition in the question style, improving the specificity of some of the symptoms, and the need for differential diagnosis. There were also comments on improving the quality of the LLM-generated queries. One expert commented on the inclusion of smallpox which has been eradicated for 40 years, but indicated the potential tie to the ongoing Monkeypox (Mpox) epidemic. Experts also indicated the need for patient images to provide more informative responses demonstrating the need for a multimodal version of the dataset.
@@
4 Discussion
This work is motivated by the need to evaluate LLMs on tropical and infectious diseases which present a potential out-of-distribution shift for LLM performance given that these tend to occur in the global south which may have less representation in datasets used for LLM training [asiedu2023globalizing, asiedu2024case]. This study confirms this distribution shift with the Gemini model achieving an accuracy of 61.5% and MedLM achieving an accuracy of 47.9% on clinical-style questions, significantly lower than reported performances on USMLE benchmarks (GPT: 90.2%, MedLM: 91.1% [saab2024capabilities]). We find that simple in-context prompting with the dataset, and potentially intentional training focus on these diseases improves the LLM performance significantly. We find that Gemini Ultra performs better than MedLM medium, however this is likely due to differences in model sizes.
Our evaluations demonstrate that including additional context such as risk factors and location in addition to symptoms also improves model performance. However indicating locations where the disease is less likely to occur, reduces model performance. This has implications for LLMs that may attempt to provide a disease diagnosis based on autorecognition of location without considering contexts such as visits to endemic locations. On this dataset, we do not find significant differences in LLM performance across race or gender. We also do not find any significant differences in performance between different styles of question asking (semantic, clinical, and consumer variations).
Our analysis reveal that LLMs tend to more accurately identify common diseases, or diseases with very specified symptoms such as Trachoma, Rabies and Yaws, while less common and less specific diseases are mislabeled. For instance we find that LLMs mostly classify Hepatitis E as Hepatitis A, though this was a common mistake among human experts as well.
Our human expert baseline finds that for both SAQs and MCQs, experts scored lower in accuracy on the full context questions than the model except in cases where we looked at expert any/at least one score. Avian influenza is an example of a disease that the LLM had trouble with, but that majority of the experts got right. On the other hand the model got diseases like Giardiasis and Onchocerciasis right, while most experts got these wrong. We do note that experts were asked not to consult any external material, which would not be the case in real-world scenarios, and real world scenarios may also provide confirmatory tests for the diseases.
Experts generally rated the dataset highly on axes of accuracy, completeness, timeliness and diversity across tropical and infectious diseases. However they suggested improvements in diversity in question asking styles, and addition of images to the questions where applicable.
Implications for Policy and Practice: Our findings demonstrate the discrepancy in LLM performance on tropical and infectious diseases, compared to reported performance on USMLE questions, identifying the need for contextual evaluation of LLMs that are used in clinical settings in the global south - i.e the need for LLM usage to take into account contextual, regional-specific and patient-specific factors. However, we also find that Gemini performs better on this dataset compared to human experts for identifying TRINDs from text-based descriptions. For healthcare workers, our findings highlight the potential of LLMs to serve as valuable decision-support tools, thus enhancing clinical diagnostic accuracy in resource-limited settings. Notably, these tools should complement, not replace, clinical judgment and should be balanced by continuous evaluation and refinement of these models to maintain their relevance and reliability in diverse clinical settings.
Limitations and Future work: Limitations of this work include the focus on only disease classification, primary focus on English and primary focus on text-based queries. Future work could explore evaluating other tasks such as management steps and treatments, additional languages and multimodal datasets such as relevant disease-related images, or sounds from coughs and breathing. Another limitation is that we used a relatively small sample size of experts, which may not represent the breath of experience in this very broad subject area. Future work should look at a larger sample size of experts reflecting geographic and sub-specialty diversity to improve expert assessment and provide a more generalizable human expert baseline.
@@ 5 Conclusion
Overall this study finds that while LLM performance on providing a diagnosis of the tropical and infectious diseases dataset is low, we find that experts performance is similarly low in most cases. LLM performance improves with simple in-context learning with our dataset. We find that larger generalist models outperform smaller specialist models. We underline the importance of context in performance, noting that providing symptoms, risk factors and location outperform the provision of symptoms alone to the LLM. This work provides a scalable methodology for evaluating LLMs for health in global settings for out-of-distribution cases.
[…]